A. Einstein mass/energy:

E = mc^{2}

B. Maxwell’s Eqations in Vacuum:

\begin{matrix}∇ · B = 0 \\ ∇ × B = \frac{1}{c} \frac{∂E}{∂t} \end{matrix} \begin{matrix}∇ · E = 0 \\ ∇ × E = −\frac{1}{c} \frac{∂B}{∂t} \end{matrix}

C. Stokes’ Theorem

∫_{M} dω = ∫_{∂M} ω

D. The boundary of a boundary is zero ∂∂ = 0

E. Heisenberg Uncertainty Principle: ΔxΔp ≥ ℏ/2

F. Primes and the zeta function:

∏_{p} \frac{1}{1 − \frac{1}{p^{s}}} = ∑_{n=1}^{∞} \frac{1}{n^{s}}

G. Kepler’s Second Law

\frac{dθ}{dt} ∝ \frac{1}{r^{2}}

H. Newton’s mass-acceleration law: F = m a I. Kepler’s Third Law

T ∝ a^{\frac{3}{2}}

J. Newton’s Law of Gravitation:

F = \frac{Gm_{1}m_{2}}{r^{2}}

K. Einstein’s General Relativity Equation:

R_{μν} − \frac{1}{2}Rg_{μν} = 8πGT_{μν}

L. Schrödinger’s Equation, Dirac Equation

iℏ\frac{∂ψ}{∂t} = −\frac{ℏ^{2}}{2m}∇^{2}ψ + V ψ

(i D̸ − m) ψ = 0

M. Atiyah-Singer Theorem for Twisted Dirac Operator:

dim ker / D_{E} − dim coker / D_{E} = ∫_{M} Â(M) · ch(E)

N. Yang-Mills equations:

F = dA + A ∧ A
D_{A}^{∗}F = 0

O.

\{\overline{Q}, Q\} = / P